# Deep Discrete Encoders: Identifiable Deep Generative Models for Rich Data with Discrete Latent Layers

### v1

@@ 7 Discussion
4000 EBIC 99.5 0.5
LRT 96.0 4.0
Spectral 2.75 97.25
6000 EBIC 98.0 2.0
LRT 93.0 7.0
Spectral 2.0 98.0
8000 EBIC 97.25 2.75
LRT 94.25 5.75
Spectral 2.0 98.0
Bernoulli, $\mathcal{B}_{g}$ 500 EBIC 0.75 99.25
@@
Parfam, value $N$ Method \ $\widehat{K}^{(2)}$ 1 2 3
Poisson, $\mathcal{B}_{s}$ 500 EBIC 99.75 0.25
LRT 99.5 0.5
Spectral 0.25 99.75
1000 EBIC 99.75 0.25
LRT 99.75 0.25
Spectral 100.0
2000 EBIC 99.25 0.75
LRT 99.0 1.0
Spectral 100.0
4000 EBIC 98.25 1.75
LRT 98.25 1.75
Spectral 100.0
6000 EBIC 99.0 1.0
LRT 98.25 1.75
Spectral 100.0
8000 EBIC 98.5 1.5
LRT 98.5 1.5
Spectral 100.0
@@
We work with the preprocessed version of the default training set, which consists of $60,000$ images, each containing information on the $28^{2}$ pixel values. Initially, each pixel takes integer values between 00 and $255$. As the data values are highly concentrated around values near 0 or larger than 200, we transform the data into binary responses by thresholding at a value of 128. In other words, pixels with values exceeding 128 are assigned a binary value of 1, while the rest are set to 0. For the sake of easier presentation and computational efficiency, we work on the subset of the dataset whose true digit labels are equal to $0,1,2,3$, and also discard some pixels with uniformly small values by selecting $J=264$ pixels whose average pixel values are larger than 40. After preprocessing, the training set consists of $N=20,679$ unlabeled images. We identically preprocess the test set, which leads to a total of $N=4,157$ images.
@@
TIMSS Response Time Data.
We use the preprocessed dataset provided in Lee and Gu, (2024). Additionally, we remove all rows of the data matrix with missing values. While our estimation procedure can be modified to handle missing data (missing at random) by modifying the likelihood function, we take this additional preprocessing step for the sake of consistency with other parts of the paper.
@@
TIMSS Response Time Data.
The dataset comes with a rough classification of the seven latent skills into $K^{(2)}=2$ categories: content skills and cognitive skills. However, we observed that fitting $K^{(2)}=2$ results in one of the proportion parameters exhibiting an unusually large value of $p_{2}=0.96$. This indicates that $A_{2}^{(2)}=1$ for $96\%$ of the students, and this skill is redundant. Hence, we have adjusted the number of higher-order latent variables to $K^{(2)}=1$, which partitions the students into two groups based on the value of $A^{(2)}_{1}$. The EBIC for the two cases were also comparable and justified this choice.
S.5.3 Performance Evaluation for the 20 Newsgroups Dataset
Perplexity.
Implementations of existing topic modeling algorithms.
S.5.4 Additional Visualization of MNIST Dataset

### v2

@@ 7 Discussion
4000 EBIC 0 99.5 0.5
LRT 0 96.0 4.0
Spectral 2.75 97.25 0
6000 EBIC 0 98.0 2.0
LRT 0 93.0 7.0
Spectral 2.0 98.0 0
8000 EBIC 0 97.25 2.75
LRT 0 94.25 5.75
Spectral 2.0 98.0 0
Bernoulli, $\mathcal{B}_{g}$ 500 EBIC 0.75 99.25 0
@@
Parfam, value NN Method \ $\widehat{K}^{(2)}$ 1 2 3
Poisson, $\mathcal{B}_{s}$ 500 EBIC 0 99.75 0.25
LRT 0 99.5 0.5
Spectral 0.25 99.75 0
1000 EBIC 0 99.75 0.25
LRT 0 99.75 0.25
Spectral 0 100.0 0
2000 EBIC 0 99.25 0.75
LRT 0 99.0 1.0
Spectral 0 100.0 0
4000 EBIC 0 98.25 1.75
LRT 0 98.25 1.75
Spectral 0 100.0 0
6000 EBIC 0 99.0 1.0
LRT 0 98.25 1.75
Spectral 0 100.0 0
8000 EBIC 0 98.5 1.5
LRT 0 98.5 1.5
Spectral 0 100.0 0
@@
We work with the preprocessed version of the default training set, which consists of $60,000$ images, each containing information on the $28^{2}$ pixel values. Initially, each pixel takes integer values between 0 and $255$. As the data values are highly concentrated around values near 0 or larger than 200, we transform the data into binary responses by thresholding at a value of 128. In other words, pixels with values exceeding 128 are assigned a binary value of 1, while the rest are set to 0. For the sake of easier presentation and computational efficiency, we work on the subset of the dataset whose true digit labels are equal to $0,1,2,3$, and also discard some pixels with uniformly small values by selecting $J=264$ pixels whose average pixel values are larger than 40. After preprocessing, the training set consists of $N=20,679$ unlabeled images. We identically preprocess the test set, which leads to a total of $N=4,157$ images.
@@
TIMSS Assessment Data.
We use the preprocessed dataset provided in Lee and Gu, (2024). Additionally, we remove all rows of the data matrix with missing entries, which resulted in a total of $N=435$ observations. While our estimation procedure can be modified to handle missing data (missing at random) by modifying the likelihood function, we take this additional preprocessing step for the sake of consistency with other parts of the paper.
This dataset includes additional information regarding the latent structure $\mathbf{G}^{(1)}$, which is the so-called $Q$-matrix in the cognitive diagnostic modeling literature (von Davier, , 2008; Lee and Gu, , 2024). The provisional $\mathbf{G}^{(1)}$ matrix specifies $K^{(1)}=7$ latent cognitive skills (Number, Algebra, Geometry, Data and Probability, Knowing, Applying, and Reasoning) as well as the skills that are required to solve each item. That is, $g^{(1)}_{j,k}=1$ if item $j$ requires latent skill $k$ to solve it. Hence, following the confirmatory latent variable modeling convention in psychometrics, we estimate the DDE parameters by fixing $\mathbf{G}^{(1)}$ to this given structure.
@@
TIMSS Assessment Data.
This dataset includes additional information regarding the latent structure $\mathbf{G}^{(1)}$, which is the so-called $Q$-matrix in the cognitive diagnostic modeling literature (von Davier, , 2008; Lee and Gu, , 2024). The provisional $\mathbf{G}^{(1)}$ matrix specifies $K^{(1)}=7$ latent cognitive skills (Number, Algebra, Geometry, Data and Probability, Knowing, Applying, and Reasoning) as well as the skills that are required to solve each item. That is, $g^{(1)}_{j,k}=1$ if item $j$ requires latent skill $k$ to solve it. Hence, following the confirmatory latent variable modeling convention in psychometrics, we estimate the DDE parameters by fixing $\mathbf{G}^{(1)}$ to this given structure.
Additionally, the dataset comes with a rough classification of the seven latent skills into $K^{(2)}=2$ categories: content skills and cognitive skills. However, we observed that fitting $K^{(2)}=2$ results in one of the proportion parameters exhibiting an unusually large value of $p_{2}=0.96$. This indicates that $A_{2}^{(2)}=1$ for $96\%$ of the students, and this skill is redundant. Hence, we have adjusted the number of higher-order latent variables to $K^{(2)}=1$, which partitions the students into two groups based on the value of $A^{(2)}_{1}$. The EBIC for the two cases were also comparable and justified this choice.
S.5.3 Additional Visualization and Performance Evaluation for the MNIST Dataset
S.5.3.1 Additional visualization
@@
S.5.3.2 Comparison with VAEs and iVAEs
We provide additional numerical comparison against VAEs and identifiable VAEs (iVAE), alongside implementation details. We use the standard VAE and iVAE in Khemakhem et al., (2020) with varying latent dimensions $K=2,5$ (these numbers are motivated by the latent dimensions $K^{(1)}=5,K^{(2)}=2$ used for DDEs). The values presented in Table 4 in the main text correspond to the VAE with latent dimension $K=2$. While we use identical neural network architectures (two-layer perceptrons with 50 hidden variables in the middle layer) and factorized priors for both VAE and iVAE, the latter requires auxiliary information to construct “label priors” while the former does not. Hence, we have provided the true digit labels as auxiliary data for iVAEs.
Accuracy 2-DDE VAE ($K=2$) VAE ($K=5$) iVAE ($K=2$) iVAE ($K=5$)
Train classif. (%) 92.0 97.1 97.5 97.5 98.2
Test classif. (%) 92.6 92.0 92.1 92.5 93.2
Train recon. (%) 79.6 82.5 83.4 82.5 86.6
Test recon. (%) 79.9 82.7 83.6 82.8 86.6
Table S.17: Performance comparison of DDEs versus VAEs and iVAEs with varying latent dimensions $K$ on the MNIST dataset. Note that the iVAEs are trained in a supervised manner, where the true labels are incorporated as auxiliary information.
We report the evaluation metrics in Table S.17. Compared to VAEs, DDEs have better test classification accuracy but lower reconstruction accuracy. We believe that the discrepancy in reconstruction accuracy results from the continuous latent variables in VAEs, which provide more detailed information. In contrast, for classification, the lack of a clear threshold in continuous latent variables seem to result in potential overfitting and lower test accuracy.
Compared to iVAEs, DDEs have lower performance in both metrics but this is not a fair comparison as iVAEs are weakly supervised (requiring the true label information), whereas DDEs are fully unsupervised. Focusing on the classification error, we note that iVAEs with $K=2$ latent dimensions have similar performance compared to DDEs. This suggests that DDEs are able to compete with benchmark iVAEs as well.
DDE [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
iVAE [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
${\boldsymbol{\beta}}_{0,0}^{(1)}$ ${\boldsymbol{\beta}}_{0,1}^{(1)}$ ${\boldsymbol{\beta}}_{0,2}^{(1)}$ ${\boldsymbol{\beta}}_{0,3}^{(1)}$ ${\boldsymbol{\beta}}_{0,4}^{(1)}$ ${\boldsymbol{\beta}}_{0,5}^{(1)}$
Table S.18: From left to right: Estimated basis images (reshaped last-layer coefficients) from the MNIST data. Brown indicates positive values and green indicates negative values. The basis images under DDEs are distinct, smoother, and more interpretable whereas that under the iVAE exhibit overlaps.
Our most interesting finding is the qualitative comparison of the basis images learned via DDEs versus iVAEs (with $K=2$) in Table 2. Here, the iVAE basis images are the last-layer perceptron coefficients. We see that the third, fourth, fifth column of the iVAE panel are almost identical. This strongly indicates that the estimated perceptron (neural network) parameters suffer from overparametrization. Consequently, it is challenging to interpret the shallow layer in the perceptron. In contrast, the basis images under DDEs are distinct, smoother, as well as highly interpretable (e.g. the basis images in the first two columns directly resemble the digits 1 and 0).
S.5.3.3 Implementation details
The VAE and iVAE was implemented by the Python codes provided in Khemakhem et al., (2020), where no auxiliary information was used for VAEs and the true digit labels were used as auxiliary information for iVAEs. The DBN is implemented similar to the earlier ablation studies in Supplement 4.2. For both models, we have used default tuning parameters and architectures provided in the original codes. To compute the classification error from latent representations, we have fitted a decision tree analogous to that for DDEs as described in the main text.
S.5.4 Performance Evaluation for the 20 Newsgroups Dataset
Perplexity.
Implementations of existing topic modeling algorithms.
S.5.5 Additional Analysis of the TIMSS Assessment Dataset
We display the estimated coefficients for each modailty in Figure S.14, focusing on the first four skills (Number, Algebra, Geometry, Data&Probability). The magnitude of the coefficients are larger for the response accuracy, indicating a larger effect of the latent variables on the responses. Additionally, we observe an interesting pattern learned from the intercept values, compared to the held-out information regarding the items (constructed-response versus multiple-choice). To be specific, the constructed-response items (indexed by the rows 2, 4, 7, 15-17, 19-20, 24-26, 29) have smaller intercept values for the left panel (response accuracy) of Figure S.14 and larger intercept values for the right panel (response time). This is analogous to the previous analysis of the same dataset in (Lee and Gu, , 2024), where each mode (response accuracy and time) were analyzed separately.
Figure S.14: Estimated coefficients for (left) response accuracy, (right) response time.
Additionally, we visualize the analog of Table 6 solely-based on the response time in Table S.19, as opposed to the analysis using both modes. Here, the same lognormal-DDE was used with identical latent dimensions. While we observe a similar trend that students enjoying math are more likely to master each skill, the discrepancy between the students enjoying/not enjoying math is lower. For instance, the discrepancy between the category ‘agree a lot’ versus ‘disagree a lot’ in Table 6 of the main paper is 0.35 versus the value of 0.14 in Table S.19. This illustrates considering both data modalities help to better learn the latent variables in a more interpretable manner.
Agree a lot 0.83 0.72 0.88 0.86 0.87
Agree a little 0.80 0.72 0.85 0.84 0.85
Disagree a little 0.76 0.74 0.85 0.84 0.84
Disagree a lot 0.69 0.67 0.76 0.75 0.78
Table S.19: Average latent variable estimate for each response category for the question “Mathematics is one of my favorite subjects”, solely based on response times.
S.6 Additional Literature Review
iVAEs.
We compare DDEs versus iVAEs and its extensions. One distinction is that iVAEs essentially have only one latent layer of random variables transformed by deterministic deep neural networks, while DDEs allow multiple latent layers. Thus, DDEs are able to model the hierarchical nature of uncertainty and are more suitable for tasks like hierarchical topic modeling (Ranganath et al., , 2015).
Also, iVAEs typically require continuous latent variables, additional auxiliary variables, and its notion of identifiability still suffers from rotational ambiguity. Most iVAEs are used to model continuous observed responses (Khemakhem et al., , 2020; Moran et al., , 2022; Kivva et al., , 2022), although there exist a few variants exclusively proposed for discrete data (binary, count) as well (Hyttinen et al., , 2022). In particular, up to our knowledge, identifiability guarantees for discrete data has been only established in Hyttinen et al., (2022) under strong parametric restrictions on the encoder. [^4] The original paper Khemakhem et al., (2020) only provided identifiability theory for continuous data, although their simulation studies suggest potential extensions to discrete data. In contrast, DDEs with discrete latent variables do not require auxiliary variables, do not suffer from rotational ambiguity, and the identifiability of DDEs holds regardless of the observed data types.
Psychometrics.